# ZIM: Zero-Shot Image Matting for Anything

### v1

@@ 6 Ablation Study
Table 4: Analysis of ZIM using box prompt evaluations: (a) Effect of Attn (prompt-aware masked attention) and Dec (hierarchical pixel decoder). (b) Effect of the masked attention in T2I (token to image) and I2T (image to token) cross-attention layers.
@@
7 Downstream Task
In this section, we demonstrate the versatility of our zero-shot image matting model, ZIM, by applying it to a variety of downstream tasks. We show that ZIM surpasses SAM in these tasks, especially in scenarios requiring higher precision in mask generation.
7.1 Zero-Shot Image Matting
To showcase the generalizability of ZIM, we evaluate it across 23 diverse datasets, including ADE20K [66], BBBC038v1 [5], Cityspcaes [12], DOORS [42], EgoHOS [65], DRAM [11], GTEA [34, 17], Hypersim [48], IBD [7], iShape [57], COCO [35], NDD20 [54], NDISPark [10, 9], OVIS [43], PIDRay [55], Plittersdorf [20], PPDLS [39], STREETS [50], TimberSeg [18], TrashCan [21], VISOR [14, 13], WoodScape [62], and Zero Waste-f [4]. Using the Automatic Mask Generation strategy introduced by SAM, we apply a regular grid of point prompts to each image and perform post-processing with thresholding and non-maximum suppression (NMS) to generate the final matting masks. Figure 10 and Figure 11 show that ZIM produces high-quality matte anything results for all datasets with considerably detailed matte quality and powerful generalization capability. Although SAM shows powerful generalization capability, the output mask is the coarse quality. In addition, existing interactive matting methods ($i.e.,$ Matte-Any, Matting-Any, and SMat) often fail to generalize the unseen data.
7.2 Image Inpainting
Image inpainting is an important application in generative AI, where precise mask generation plays a critical role in removing or reconstructing parts of an image. We follow the Inpaint Anything framework [64], allowing users to interactively select objects to be erased by clicking on them. In this experiment, we compare the performance of SAM and ZIM on the COCO 2017 validation set [35]. To generate object masks, we randomly sample 10 points from the COCO ground-truth segmentation masks and use them as point prompts for both models. The resulting masks are then passed to an inpainting model. As shown in Figure 6, ZIM produces more accurate object masks than SAM, leading to significantly better inpainting results. For complex objects like flowerpots and hair, SAM’s coarse masks fail to remove the object cleanly, leaving noticeable artifacts. In contrast, our precise matting masks enable the inpainting model to smoothly remove objects without artifacts.
Furthermore, to quantify this performance, we use CLIP Distance and CLIP Accuracy [16, 61] as evaluation metrics. CLIP Distance measures the similarity between the source and inpainted regions, where a larger distance indicates better removal. CLIP Accuracy evaluates the change in class predictions after removal, considering the task successful if the original class is absent from Top-1/3/5 predictions. Following [61], we use the text prompt a photo of a {category name}. Table 5 shows that ZIM outperforms SAM on both metrics by better preserving surrounding context after inpainting through enhanced mask quality.
7.3 3D Object Segmentation with NeRF
The quality of the segmentation mask is crucial when converting a 2D mask into a 3D representation. In this work, we adopt the SA3D framework [6], which utilizes SAM to segment 3D objects from 2D masks by manually prompting the target object in a single view. By replacing SAM with ZIM in the 2D mask segmentation process, we significantly improve the quality of the resulting 3D objects. Since the SA3D framework relies on binary masks for projecting 2D masks into 3D space, we binarized the ZIM results using a threshold of 0.3. For qualitative evaluation, we use the LLFF [38] and 360 [3] datasets, which contain fine-grained object details. For quantitative evaluation, we employ the NVOS [47] dataset, which includes finely annotated 2D masks. We fixed the number of self-prompting points at 10 and followed the experimental setup of SA3D [6] for pre-trained NeRFs and manual prompts.
Figure 7 presents the qualitative results of segmented 3D objects guided by the SAM and ZIM models on the LLFF-trex, LLFF-horns [38], and 360-kitchen (Lego) [3] datasets. Compared to SAM, which often misses finer details due to its coarse-level mask generation, ZIM captures more intricate object features. Furthermore, as shown in Table 6, ZIM outperforms SAM in quantitative 3D segmentation results on the target view, reporting higher mask IoU scores. These findings demonstrate that the precise matting capabilities of ZIM extend beyond 2D tasks, significantly enhancing the quality of 3D object segmentation.
7.4 Medical Image Segmentation
8 Conclusion
In this paper, we introduced ZIM, a pioneering zero-shot image matting model designed to address the limitations of SAM and existing zero-shot matting approaches in generating high-quality matte masks. While SAM’s extensive training on the SA1B dataset enables broad generalization across segmentation tasks, it lacks the mask precision needed for tasks like image matting. Existing matting approaches that fine-tune SAM on public matting datasets achieve higher precision but often sacrifice zero-shot versatility due to reliance on macro-level labels. To overcome these challenges, we developed ZIM by constructing the SA1B-Matte dataset through a novel label conversion method, allowing for micro-level matting without extensive manual annotation. Supported by strategies like Spatial Generalization Augmentation and Selective Transformation Learning, ZIM achieves high-fidelity matte labels while preserving SAM’s zero-shot potential. Additionally, we enhanced ZIM with a hierarchical pixel decoder and prompt-aware masked attention mechanism, allowing it to produce robust, detailed masks that mitigate checkerboard artifacts and accurately respond to visual prompts.
Experimental results on the MicroMat-3K dataset confirm ZIM’s effectiveness in producing high-quality matting results in a zero-shot setting, consistently outperforming SAM and other matting models on micro-level matting tasks. Furthermore, evaluations across downstream tasks, including image inpainting, 3D NeRF, and medical image analysis, demonstrate ZIM’s versatility and applicability in fields requiring precise mask generation. In each task, ZIM not only preserves context and fine details but also improves over SAM and existing zero-shot matting approaches by delivering high-resolution masks that better capture intricate structures. We believe ZIM’s strong performance in zero-shot matting marks an important step forward in interactive image matting and inspires further exploration in developing more refined zero-shot solutions. Future work could explore ZIM’s adaptability in additional fields such as video or 3D domains, potentially extending its impact across both established and emerging areas in visual understanding.
@@ Appendix A MicroMat-3K Details
Method Backbone Latency Prompt Fine-$\downarrow$ Coarse-$\downarrow$ (ms)$\downarrow$ MSE MSES MSEM MSEL MSE MSES MSEM MSEL SAM [27] ViT-B [15] 172.5 point 21.651 2.717 7.145 70.502 5.569 4.092 5.405 77.761 box 11.057 0.329 2.983 38.501 1.044 0.181 2.456 24.519 ViT-L [15] 361.6 point 15.663 3.312 5.165 49.202 4.293 3.640 3.389 46.278 box 7.989 0.320 2.423 27.276 0.534 0.145 1.606 5.575 ViT-H [15] 605.2 point 14.534 3.619 5.753 43.371 2.100 0.653 3.278 56.086 box 6.188 0.281 1.687 21.389 0.468 0.152 1.334 4.610 SAM2 [45] Hiera-B+ [49] 147.4 point 25.296 15.657 15.970 53.346 14.794 14.786 12.128 49.058 box 12.024 0.322 2.155 43.670 0.613 0.152 1.600 10.194 Hiera-L [49] 195.0 point 16.937 0.871 5.613 56.702 2.572 0.954 4.308 57.804 box 10.616 0.263 1.888 38.636 0.704 0.149 1.424 18.019 HQ-SAM [23] ViT-B [15] 177.2 point 36.674 5.094 9.356 123.199 6.457 3.602 9.596 102.834 box 42.457 0.392 4.369 160.456 2.733 0.230 2.521 124.372 ViT-L [15] 368.1 point 20.481 1.928 6.496 67.979 3.046 1.200 4.987 66.373 box 19.881 0.326 2.683 73.913 0.762 0.163 1.408 21.129 ViT-H [15] 608.1 point 22.547 5.308 6.601 71.446 3.599 1.659 5.034 77.643 box 23.743 0.263 2.538 89.518 0.789 0.164 1.266 24.469 Matte-Any [59] ViT-B [15] 668.5 point 20.844 3.381 7.953 65.108 6.053 4.739 5.897 70.443 box 9.746 0.918 3.856 31.109 1.983 1.235 3.039 24.426 ViT-L [15] 814.1 point 15.230 3.985 6.243 44.842 5.116 4.570 3.996 44.606 box 7.323 0.975 3.692 21.711 1.597 1.222 2.418 9.119 ViT-H [15] 1036.9 point 14.119 4.270 7.085 38.704 3.022 1.669 3.913 56.044 box 6.048 0.917 2.992 17.872 1.571 1.228 2.294 8.836 Matting-Any [33] ViT-B [15] 200.3 point 77.335 27.256 41.349 202.692 36.187 31.549 40.993 197.340 box 68.372 18.286 33.111 192.566 23.780 17.344 36.194 175.418 SMat [60] ViT-B [15] 263.4 point 123.664 31.549 40.993 197.340 59.113 52.305 68.709 250.587 box 133.515 17.344 36.194 175.418 61.157 53.280 74.099 261.544 ZIM (ours) ViT-B [15] 187.8 point 8.213 0.870 3.962 24.934 1.788 1.444 1.731 18.983 box 1.893 0.205 2.228 3.617 0.448 0.200 1.382 0.632 ViT-L [15] 373.4 point 5.825 0.563 2.724 17.898 1.719 1.041 1.574 35.687 box 1.589 0.191 1.888 2.982 0.446 0.175 1.458 0.686
@@
Appendix B Detailed Experiemental Result
Table 7 provides an in-depth evaluation of zero-shot matting models, with additional experiments utilizing various backbone networks: ViT-L and ViT-H [15] for SAM [27], HQ-SAM [23], and Matte-Any [59], and Hiera-L [49] for SAM2 [45]. Furthermore, we investigate model behavior based on the size of the target object by defining three object size groups, according to the ratio of the foreground region in the image: small (ratio $<$ 1%), medium (1% $\leq$ ratio $<$ 10%), and large (ratio $\geq$ 10%). The MSE error is reported for each object size group, offering a detailed understanding of model performance across varying object sizes. Additionally, we measure the model throughput using an NVIDIA V100 GPU to assess computational efficiency.
@@ Appendix B Detailed Experiemental Result
Figure 13: Qualitative samples of text prompting results on the RefMatte-RW100 [32] dataset.
Appendix C Expanding Prompt Sources
Interactive models, such as SAM, commonly support only point and box prompts. Here, we demonstrate the potential ZIM offers a more flexible approach by expanding the variety of prompt types, including text and scribble prompts.
@@ Appendix C Expanding Prompt Sources
To enable text prompts, we integrate ZIM with the Grounded-SAM framework [46]. Grounded-SAM uses a grounding object detection model that processes an image-text pair and returns bounding boxes for objects mentioned in the text. ZIM then uses these bounding boxes as prompts to produce detailed matte outputs. We refer to this combined model as Grounded-ZIM. As shown in Figure 13, Grounded-ZIM provides high-quality outputs with a simple text prompting pipeline, offering more precise and robust mask generation than Grounded-SAM.
@@
In addition, ZIM can support scribble prompts, which provide users with an intuitive way to mark regions of interest. We implement this functionality by sampling points along the scribble path. To ensure comprehensive coverage of the scribble region, we employ uniform sampling with setting the maximum number of sampled points to 24. It allows ZIM to effectively handle the scribble input and generate high-quality matte outputs. Figure 14 shows an example of how the scribble prompt leads to accurate and stable results. These expansions highlight the versatility of ZIM in accommodating diverse input prompts.
Figure 14: Qualitative examples of scribble prompting results.
Appendix D Additional Ablation Study
In this section, we explore the impact of hyperparameters on the performance of the ZIM model, focusing on $\sigma$ and $\lambda$.
@@ Appendix D Additional Ablation Study
Table 8: Analysis of ZIM using point prompt evaluations: (a) Effect of the hyperpameter $\sigma$. (b) Effect of the hyperpameter $\lambda$.
Appendix E Additional Qualitative Samples

### v2

@@ 6 Ablation Study
[…]
Table 3: Analysis of ZIM using box prompt evaluations: (a) Effect of Attn (prompt-aware masked attention) and Dec (hierarchical pixel decoder). (b) Effect of the masked attention in T2I (token to image) and I2T (image to token) cross-attention layers. (c) Effect of trainset: our SA1B-Matte and public matting datasets.
@@
Analysis of Training Dataset. Table 3(c) investigates the influence of the training dataset on ZIM’s zero-shot matting performance. When ZIM is trained on publicly available matting datasets, which predominantly contain macro-level masks, the performance on the micro-level MicroMat3K dataset significantly degrades. This result underscores the necessity of training with micro-level mask annotations for effective generalization to multi-granularity details. Despite Matting-Any [41] also being trained on SA1B-Matte, ZIM still surpasses its performance by a considerable margin, highlighting the advancements of our network architectural improvements for the zero-shot interactive matting task.
Figure 7: Domain shift between traditional public matting testsets and ZIM trained with SA1B-Matte (object-level).
Figure 8: Quantitative comparison of ZIM (with varying numbers of point prompts) and existing matting methods (Matte-Any [79], Matting-Any [41], GFM [39], AIM [37], and SMat [80]) on the traditional matting test sets (AIM-500 [37] and P3M-500-NP [36]). The “auto” implies prompt-free mode.
Discussion on Domain Shift. There is a distinct domain shift between traditional matting test sets and ZIM trained with SA1B-Matte. Traditional matting test sets [37, 36] typically regard entire salient objects as foreground, whereas ZIM mainly focuses on object- and part-level matting (Figure 7). This mismatch leads to substantial performance penalties for ZIM under box prompting on these test sets (Figure 7), which is inherent from SAM’s prompt ambiguity issue. However, by leveraging dense multiple-point prompting, ZIM effectively mitigates this discrepancy, surpassing existing methods [79, 41, 39, 37, 80], even some of which are explicitly trained on datasets similar to the evaluation domain (Figure 8). This highlights the adaptability of our zero-shot interactive matting modeling.
7 Conclusion, Limitation, and Future Work
In this paper, we presented a pioneering zero-shot image matting model that advances the field by generating precise, fine-grained matte masks. We addressed the limitations of SAM, which struggles with high-detail segmentation tasks, by introducing a novel label conversion method and enhancing the network architecture with a hierarchical pixel decoder and prompt-aware masked attention mechanism. However, ZIM inherits inherent shortcomings from SAM, including ambiguous visual prompt handling and robustness in uncertain predictions. Future research should aim at resolving these distinct challenges through innovative approaches to prompt design and uncertainty-based modeling. We hope that the research community will continue to build on this work, exploring new applications in computer vision and enhancing its zero-shot matting performance.
@@ Appendix A MicroMat-3K Details
Appendix B Downstream Task
B.1 Zero-Shot Image Matting
To showcase the generalizability of ZIM, we evaluate it across 23 diverse datasets, including ADE20K [87], BBBC038v1 [7], Cityspcaes [16], DOORS [56], EgoHOS [86], DRAM [15], GTEA [43, 22], Hypersim [63], IBD [9], iShape [76], COCO [45], NDD20 [71], NDISPark [14, 13], OVIS [57], PIDRay [72], Plittersdorf [26], PPDLS [52], STREETS [65], TimberSeg [23], TrashCan [27], VISOR [18, 17], WoodScape [82], and Zero Waste-f [5]. Using the Automatic Mask Generation strategy introduced by SAM [34], we apply a regular grid of point prompts to each image and perform post-processing with thresholding and non-maximum suppression (NMS) to generate the final matting masks. Figure 10 and Figure 11 show that ZIM produces high-quality matte anything results for all datasets with considerably detailed matte quality and powerful generalization capability. Although SAM shows powerful generalization capability, the output mask is the coarse quality. In addition, existing interactive matting methods ($i.e.,$ Matte-Any [79], Matting-Any [41], and SMat [80]) often fail to generalize the unseen data.
B.2 Image Inpainting
Image inpainting is an important application in generative AI, where precise mask generation plays a critical role in removing or reconstructing parts of an image. Following the Inpaint Anything framework [84], we guide SAM and ZIM masks to the inpainting model [70]. As shown in Figure 12, ZIM produces more accurate object masks than SAM, leading to significantly better inpainting results. For complex objects like flowerpots and hair, SAM’s coarse masks fail to remove the object cleanly, leaving noticeable artifacts. In contrast, our precise matting masks enable the inpainting model to smoothly remove objects without artifacts. To provide quantitative analysis, we use CLIP Distance and CLIP Accuracy [21, 81] as evaluation metrics. CLIP Distance measures the similarity between the source and inpainted regions, where a larger distance indicates better removal. CLIP Accuracy evaluates the change in class predictions after removal, considering the task successful if the original class is absent from Top-1/3/5 predictions. Following [81], we use the text prompt a photo of a {category name}. Table 4 shows that ZIM outperforms SAM on both metrics by better preserving surrounding context after inpainting through enhanced mask quality.
B.3 3D Object Segmentation with NeRF
The quality of the segmentation mask is crucial when converting a 2D mask into a 3D representation. In this work, we adopt the SA3D framework [8], which utilizes SAM to segment 3D objects from 2D masks by manually prompting the target object in a single view. By replacing SAM with ZIM in the 2D mask segmentation process, we significantly improve the quality of the resulting 3D objects. Figure 13 presents the qualitative results of segmented 3D objects guided by the SAM and ZIM models on the LLFF-trex and LLFF-horns [51] dataset. Compared to SAM, which often misses finer details due to its coarse-level mask generation, ZIM captures more intricate object features. These findings demonstrate that the precise matting capabilities of ZIM extend beyond 2D tasks, significantly enhancing the quality of 3D object segmentation. For quantitative evaluation, we employ the NVOS [62] dataset, which includes finely annotated 2D masks. Since the SA3D framework relies on binary masks for projecting 2D masks into 3D space, we binarized the ZIM results using a threshold of 0.3. We fixed the number of self-prompting points at 10 and followed the experimental setup of SA3D [8] for pre-trained NeRFs and manual prompts. As shown in Table 5, ZIM outperforms SAM in quantitative 3D segmentation results on the target view, reporting higher mask IoU scores. These findings demonstrate that the precise matting capabilities of ZIM extend beyond 2D tasks, significantly enhancing the quality of 3D object segmentation.
B.4 Medical Image Segmentation
[…]
@@ B.4 Medical Image Segmentation
Appendix C Additional Experiemental Results
C.1 Detailed Experiment Settings
We evaluate interactive matting performance using point and box prompts derived from ground-truth masks. For point sampling, we adopt RITM [66] with a maximum of 12 points. Box prompts are extracted from the min-max foreground coordinates, with up to 10% random perturbation to assess robustness to noisy box inputs. This sampling strategy is applied to MicroMat3K and public matting datasets such as AIM-500 [37] and P3M-500-NP [36].
ZIM is trained on 1% of SA1B-Matte, containing 2.2M matte labels. Since SAM pre-trained weights are utilized, increasing the training data to 10% did not yield noticeable improvements, indicating that 1% suffices for enhancing fine-grained representations.
All experiments are conducted using 8 V100 GPUs. Training the label converter and ZIM takes approximately 7 and 5 days, respectively.
C.2 Detailed Quantitative Comparison
Table 6 provides an in-depth evaluation of zero-shot matting models, with additional experiments utilizing various backbone networks: ViT-L and ViT-H [20] for SAM [34], HQ-SAM [29], and Matte-Any [79], and Hiera-L [64] for SAM2 [60]. Furthermore, we investigate model behavior based on the size of the target object by defining three object size groups, according to the ratio of the foreground region in the image: small (ratio << 1%), medium (1% $\leq$ ratio << 10%), and large (ratio $\geq$ 10%). The MSE error is reported for each object size group, offering a detailed understanding of model performance across varying object sizes. Additionally, we measure the model throughput using an NVIDIA V100 GPU to assess computational efficiency.
@@ C.2 Detailed Quantitative Comparison
C.3 Discussion on Transparency Prediction
In the image matting task, there is a distinct scenario, that is, predicting the transparency of objects, such as glasses and fire. Due to its inherent difficulties, existing matting methods [49, 83, 6] typically rely on curated transparency datasets [6] containing close-up transparent object images with clear backgrounds, making them effective in constrained environments but less adaptable to open-world scenarios. ZIM, on the other hand, is designed for fine-grained mask representation in general object segmentation across diverse and complex scenes. However, predicting transparency within ZIM remains a challenge due to the lack of large-scale open-world transparency datasets and the inherent complexity of identifying transparent regions in natural images.
Model Input Transparent-460 [6] MSE $\downarrow$ SAD $\downarrow$ IndexNet [49] Trimap 112.53 573.09 MGMatting [83] Trimap 6.33 111.92 TransMatting [6] Trimap 4.02 88.34 ZIM (ours) Box 15.55 298.78
Table 7: Quantitative results on the transparent object matting dataset, Transprent-460 [6].
To investigate ZIM’s adaptability to transparency prediction, we fine-tune it on the Transparent-460 dataset [6] and compare its performance against specialized transparency matting methods [49, 83, 6], as shown in Table 7. While these methods leverage trimaps to provide explicit spatial guidance for transparency estimation, ZIM relies only on sparse box prompts, which offer less precise object boundary information. Despite this limitation, ZIM achieves reasonable results, demonstrating its strong transferability.
C.4 Expanding Prompt Sources
Interactive models, such as SAM, commonly support only point and box prompts, which can be limiting for complex scenarios requiring more intuitive user interaction. Here, we demonstrate that ZIM offers a more flexible approach by expanding the variety of prompt types, including text and scribble prompts, enabling more versatile and user-friendly matting applications.
Figure 16: Qualitative samples of text prompting results using Grounded-ZIM on the RefMatte-RW100 [40] dataset. Grounded-ZIM replaces SAM with ZIM in the Grounded-SAM framework [61], leveraging Grounding DINO [46]’s open-set object detection with ZIM’s superior matting capabilities to produce more precise and robust results than the original Grounded-SAM.
@@ C.4 Expanding Prompt Sources
To enable text prompts, we integrate ZIM with the Grounded-SAM framework [61]. Grounded-SAM combines Grounding DINO [46], an open-set object detector, with the Segment Anything Model (SAM) to process image-text pairs and return bounding boxes for objects mentioned in the text. In our approach, we replace SAM with ZIM, constructing Grounded-ZIM that leverages Grounding DINO’s robust detection capabilities while utilizing ZIM’s superior matting performance for fine-grained boundary refinement. This combination enables natural language-driven matting without requiring precise manual annotation. ZIM then uses the detected bounding boxes as prompts to produce detailed matte outputs with enhanced edge quality and alpha channel precision. As shown in Figure 16, Grounded-ZIM provides high-quality outputs with a simple text prompting pipeline, offering more precise and robust mask generation than the original Grounded-SAM framework.
@@
In addition, ZIM can support scribble prompts, which provide users with an intuitive way to mark regions of interest without requiring precise boundary delineation. We implement this functionality by sampling points along the scribble path, converting the continuous stroke into discrete point prompts that ZIM can process effectively. To ensure comprehensive coverage of the scribble region while maintaining computational efficiency, we employ uniform sampling with the maximum number of sampled points set to 24. This design strikes a balance between prompt coverage and processing speed, allowing ZIM to effectively handle the scribble input and generate high-quality matte outputs with smooth boundaries. Figure 17 shows an example of how the scribble prompt leads to accurate and stable results across diverse scenarios. These expansions highlight the versatility of ZIM in accommodating diverse input modalities, making it suitable for various real-world applications requiring different levels of user interaction.
Figure 17: Qualitative examples of scribble prompting results. ZIM processes scribble inputs through uniform point sampling (max 24 points) along the scribble path, generating accurate and stable matte outputs for marked regions of interest.
C.5 Additional Ablation Study
Impact of our proposed components on existing segmentation models.
Table 8 demonstrates that both the Label Converter and SA1B-Matte training data provide substantial improvements to baseline performance on the MicroMat3K fine-grained dataset. The Label Converter, when applied to SAM’s coarse output masks, reduces MSE from 11.05 to 7.34, representing a significant improvement in fine-grained mask quality. However, the converter cannot address inherent limitations in SAM’s base predictions, such as checkerboard artifacts, resulting in performance that remains below ZIM’s 1.89 MSE.
In addition, training existing models ($i.e.,$ SAM, HQ-SAM, and Matting-Any) on our SA1B-Matte dataset yields consistent improvements across all models. SAM trained with SA1B-Matte achieves 2.71 MSE, while HQSAM and Matting-Any reach 13.35 and 12.26 MSE, respectively. Despite these improvements, all existing models underperform compared to ZIM, highlighting the importance of our architectural innovations, including the Hierarchical Pixel Decoder and Prompt-Aware Masked Attention. The performance gap is particularly notable for HQ-SAM and Matting-Any, which rely on the frozen SAM model that limits their capacity to learn fine-grained representations during training.
[…]
Table 8: Analysis of our proposed components ($i.e.,$ Label Converter and SA1B-Matte dataset) when applied to existing models.
@@ C.6 Additional Qualitative Samples
Table 9: Analysis of ZIM using point prompt evaluations: (a) Effect of the hyperparameter $\sigma$ and (b) Effect of the hyperparameter $\lambda$.